You want to buy a house that costs $280,000. You will make a down payment equal to 15 percent of the price of the house and finance the remainder with a loan that has an interest rate of 5.47 percent compounded monthly. If the loan is for 25 years, what are your monthly mortgage payments?
Answer #1

Loan Amount = House Value * (1 - Downpayment%)

Loan Amount = $280,000 * (1 - 15%) = $238,000

Then we use PV of annuity formula, according to which

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

n = 25 * 12 = 300 months

r = 5.47%/12 = 0.4558% (monthly)

238000 = P* 1-(1+0.004558)-300 0.004558

238000 = P * 163.3194

P = $1,457.27